OpenStudy (precal):

Calculus Population y grows according to the equation dy/dt=ky, where k is a constant and t is measured in years. If the population doubles every 10 years, then the value of k is ?

OpenStudy (anonymous):

You'll have to integrate \[\frac{ dy }{ dt } = k y \implies \frac{ dy }{ y } = k dt \implies \int\limits \frac{ dy }{ y } = \int\limits k dt\]

OpenStudy (anonymous):

(Yes, it's a differential equation)

OpenStudy (precal):

ln absolute value (y)=kt+c

OpenStudy (precal):

then I use e as a base for both side

OpenStudy (anonymous):

Yup :)

OpenStudy (precal):

y=2e^(kt)

OpenStudy (precal):

I meant y=Ce^(kt)

OpenStudy (precal):

not sure what to do afte that

OpenStudy (precal):

I meant after that

OpenStudy (anonymous):

Well if you remember from earlier population and growth problems, the initial population is c0 at t = 0 (initial conditions). so we have y = c0 which implies ln c0 = 0+C hence our constant C = lnc0 \[lny = kt + \ln c_0\] now we look at when the population doubles every 10 years. Can you figure that out/ finish it off?

OpenStudy (precal):

I am not sure I can finish this one on my own......

OpenStudy (anonymous):

No worries, all we have to note is that t will be 10, y = 2 c0 as it's doubling, it can be tricky at first, these problems do require lots of practice :P

OpenStudy (anonymous):

You can plug that all into our equation above, and then it's just logarithms and algebra. \[\ln(2 c_0) = 10k+\ln( c_0)\]

OpenStudy (anonymous):

Solve for k and that's your answer, I hope that made some sense :)

OpenStudy (precal):

thanks

OpenStudy (precal):

.069 is the solution and I have it on my choices to pick from Thanks alot

OpenStudy (anonymous):

Your welcome :)